Any help would be really appreciated please. Ta
This is the last of my problems for now, I am really stuck. I do not know how to get the x and the y onto the same side with the = at the end. Ta
I have solve for X and Y
y=9-3x
x=8-2y

Answer by checkley75(3666)

Y=9-3X JUST SUBSTITUTE (9-3X) FOR Y IN THE OTHER EQUATION 7 SOLVE FOR X.
X=8-2(9-3X)
X=8-18+6X
X-6X=-10
-5X=-10
X=-10/-5
X=2 ANSWER. NOW REPLACE 2 FOR X & SOLVE FOR Y
Y=9-3*2
Y=9-6
Y=3 ANSWER.
PROOF
2=8-2*3
2=8-6
2=2